When a charged particle moves at an angle of 27.2 degrees with respect to a magnetic field, it experiences a magnetic force of magnitude F. At what angle (less than 90 degrees) with respect to this field will this particle, moving at the same speed, experience a magnetic force of magnitude 2F?
Answer #1

force due to magnetic field =

here q = charge of the particle

= velocity of the particle

= magnetic field

F = qvB sin ()

so F2/F1 = sin (2)/sin (1)

here 1 = 27.20

2 = ?

F2/F1 = 2

solving

sin 2 = 66.10
Answer #2

Calculation:

The magnetic force on a charged particle moving at an angle θ to a magnetic field is given by:

F=qvBsinθ

We are given:

  1. At θ₁ = 27.2°, the force is F.

  2. We need to find θ₂ where the force is 2F, with the same qv, and B.

Set up the ratio of the forces:

2FF=qvBsinθ2qvBsinθ1

Simplify:

2=sinθ2sin(27.2°)

Solve for sinθ2:

sinθ2=2sin(27.2°)sinθ2=20.457sinθ2=0.914

Find θ₂:

θ2=sin1(0.914)θ261.6°

Thus, the angle is 61.6 degrees


answered by: Harshwardhan kunal